OpenStudy (anonymous):

Does anyone know the standard eaquation for a ellipse, hyperbola, circle ,and parabola?

OpenStudy (anonymous):

General equation of the circle centered at (a,b) with a radius r:\[(x-a)^2+(y-b)^2 = r^2\] Ellipse: \[\frac{(x-a)^2}{c^2} + \frac{(y-b)^2}{d^2} = 1\] Hyperbola:\[\frac{(x-a)^2}{c^2}-\frac{(y-b)^2}{d^2} = 1\] Parabola: \[y = ax^2 + bx + c\]